OpenStudy (anonymous):

how do you take the derivative of e^-3x

OpenStudy (anonymous):

-3e^(-3x)

OpenStudy (anonymous):

(e^x)' = e^x

OpenStudy (anonymous):

e^-3x Let u=-3x so that de^u=e^u(du), therefore de^-3x=(e^-3x)d(-3x) =(e^-3x)(-3) =-3e^-3x

OpenStudy (unklerhaukus):

if \[y = e^{ax}\]\[y' =ae^{ax}\] in your example \[a=-3\]